OpenStudy (anonymous):

write cos(-240 degrees) as a function of positive acute angles. Please explain how you got to that answer as well :) I'm just not getting how to do this. :(

OpenStudy (cwrw238):

cos(-240) = cos (360-2400 = cos 120 now use the double angle formula cos 2A = cos^2 A - sin^ A thus cos (-240) = cos120 = cos^2 60 - sin^2 60